How many proper subset do the following sets have?
a) A={7,11}
b) B={a,b,c,d}

Answer by Edwin McCravy(20055)

There are 2n subsets and 2n-1 PROPER subsets,
where n is the number of elements. The empty sub

The only IMPROPER subset is the set itself.

How many proper subset do the following sets have? 
 a) A={7,11} 

A has 2 elements so there are 22-1 = 4-1 = 3
PROPER substers.  They are { }, {7} and {11}.   

----------------

 b) B={a,b,c,d}

B has 4 elements so there are 24-1 = 16-1 = 15
PROPER substers.

They are:

1. { }
2. {a}
3. {b}
4. {c}
5. {d}
6. {a,b}
7. {a,c}
8. {a,d}
9. {b,c}
10.{b,d}
11. {c,d}
12. {a,b,c}
13. {a,b,d}
14. {a,c,d}
15. {b,c,d} 

Edwin